# Universal Incremental Learning: Mitigating Confusion from Inter- and Intra-task Distribution Randomness

### v1

@@ Abstract
Figure 1: Comparison of existing IL scenarios and our UIL scenario on a 6-class dataset with 4 domains. The horizontal axis represents class variation, while the vertical axis represents domain variation. Each incremental task is framed by a black box. UIL is a more general IL scenario than existing IL scenarios by simultaneously considering randomness in both the type and scale of increments, as the model neither knows which new classes or domains will increase across all task distributions, nor the number of new classes or domains within each task distribution.
@@ 4.2 Mitigating Confusion from Intra-task Distribution Randomness
To address this, we introduce a magnitude recalibration module to alleviate the asymmetrical optimization by rescaling the classifier’s gradient magnitude. Specifically, denote $\mathcal{G}_{m}=\{||g^{1}||,||g^{2}||,\cdots,||g^{c}||\}$ as the set of gradient magnitude, where $c$ is the class index of $c$-th weight vector. $g^{c}$ is defined in Eq. 7:
$||g^{c}||=||g^{c}_{ce}||+\gamma\cdot||g_{em}^{c}||,$ (7)
where $||g^{c}_{ce}||$ , $||g_{em}^{c}||$ and $||g^{c}||$ are gradient magnitude from $\mathcal{L}_{ce}$, $\mathcal{L}_{em}$, and total loss $\mathcal{L}$, respectively. $\gamma$ is the same balanced coefficient as Eq.5. We rescale each gradient magnitude in $\mathcal{G}_{m}$ using class-specific factor ${w}$ derived from Min-Max Normalization on $\mathcal{G}_{m}$. As shown in Eq. 8:
$||\hat{g}^{i}||$ $=w_{i}*||g^{i}||,$ (8)
where $\text{where }w_{i}$ $=\frac{||g^{i}||-\min(\mathcal{G}_{m})}{\max(\mathcal{G}_{m})-\min(\mathcal{G}_{m})}.$
@@ 5.1 Experiment Setup
Comparison Methods and Implementation Details. We compare MiCo with the state-of-the-art IL methods of three categories: prompt-based methods [39, 38, 33, 37], regularization-based methods [26, 8], mixture-based methods [7]. In addition, we also conduct the fine-tuning and joint-training experiments to provide the lower bound and upper bound of performance in a linear probe way [28, 9, 10]. All IL methods uniformly use ViT-B/16 [4] as the pre-trained backbone. Each method is trained over three random experiments with different seeds, and we report final performance as the mean and standard deviation. For training configuration (i.e., learning rate, epochs), we keep the same setting as ICON [26].
Table 1: Task configuration on three benchmarks in the UIL and VIL scenario. $T,\mathcal{C}_{t}$ and $\mathcal{M}_{t}$ denote the total number of sequential tasks, the number of classes and the number of domains within each incremental task, respectively. In the UIL scenario, $unk$ indicates that the number of classes and domains are unknown due to random scale of these increments. The value with asterisk (*) indicates that we try to ensure that each task covers multiple domains, preventing the UIL scenario from degrading into the VIL scenario for iDigits with fewer classes and domains, which leads to fewer tasks overall.
@@
$T$ $\mathcal{C}_{t}$ $\mathcal{M}_{t}$ $T$ $\mathcal{C}_{t}$ $\mathcal{M}_{t}$

### v2

@@ Abstract
Figure 1: Comparison of existing IL scenarios and our UIL scenario on a 6-class dataset with 4 domains. The horizontal axis denotes class index and the vertical axis denotes domain index. Each incremental task is outlined in black. CIL, DIL, and VIL assume that each incoming task introduces only class increments, domain increments, or both class and domain increments, respectively. In contrast, UIL is a more general IL scenario than existing ones by simultaneously introducing randomness in both the type and scale of increments, as the model neither knows which new classes or domains will increase across all task distributions, nor the number of new classes or domains within each task distribution.
@@ 4.2 Mitigating Confusion from Intra-task Distribution Randomness
To address this, we introduce a magnitude recalibration module to alleviate the asymmetrical optimization by rescaling the classifier’s gradient magnitude. Specifically, denote $\mathcal{G}_{m}=\{||g^{1}||,||g^{2}||,\cdots,||g^{c}||\}$ as the set of gradient magnitude, where $c$ is the class index of $c$-th weight vector. $||g^{c}||$ is defined in Eq. 7:
$||g^{c}||=||g^{c}_{ce}+\gamma\cdot g_{em}^{c}||,$ (7)
where $g^{c}_{ce}$ and $g_{em}^{c}$ are gradient vector from $\mathcal{L}_{ce}$ and $\mathcal{L}_{em}$, respectively. $||g^{c}||$ is the accumulated gradient magnitude after vector summation. $\gamma$ is the same balanced coefficient as Eq.5. We rescale each gradient magnitude in $\mathcal{G}_{m}$ using class-specific factor ${w}$ derived from Min-Max Normalization on $\mathcal{G}_{m}$. As shown in Eq. 8:
$||\hat{g}^{c}||$ $=w_{c}*||g^{c}||,$ (8)
where $\text{where }w_{c}$ $=1-\frac{||g^{c}||-\min(\mathcal{G}_{m})}{\max(\mathcal{G}_{m})-\min(\mathcal{G}_{m})}.$
@@ 5.1 Experiment Setup
Comparison Methods and Implementation Details. We compare MiCo with the state-of-the-art IL methods of three categories: prompt-based methods [39, 38, 33, 37], regularization-based methods [26, 8], mixture-based methods [7]. In addition, we also conduct the fine-tuning and joint-training experiments to provide the lower bound and upper bound of performance in a linear probe way [28, 9, 10]. All IL methods uniformly use ViT-B/16 [4] as the pre-trained backbone. Each method is trained over three random experiments with different seeds, and we report final performance as the mean and standard deviation. For training configuration (i.e., learning rate, epochs and learnable adapter for fine-tuning backbone), we keep the same setting as ICON [26].
Table 1: Task configuration on three benchmarks in the UIL and VIL scenario. $T$,$\left\|\mathcal{C}_{t}\right\|$ and $\left\|\mathcal{M}_{t}\right\|$ denote the total number of sequential tasks, the number of classes and the number of domains within each incremental task, respectively. In the UIL scenario, $unk$ indicates that the number of classes and domains are unknown due to random scale of these increments. The value with asterisk (*) indicates that we try to ensure that each task covers multiple domains, preventing the UIL scenario from degrading into the VIL scenario for iDigits with fewer classes and domains, which leads to fewer tasks overall.
@@
$T$ $\left\|\mathcal{C}_{t}\right\|$ $\left\|\mathcal{M}_{t}\right\|$ $T$ $\left\|\mathcal{C}_{t}\right\|$ $\left\|\mathcal{M}_{t}\right\|$